Corrections

* The follwoing page numbers are those in the Journal. page 72 line -3 The space defined in line −1 does not correspond to the Besov space when β ∈ Z. We thus assume β ∉ Z here. page 74 line 9 From Theorem 2.7 to Theorem 2.11, we should change the definition of D(E), Ẽ as

D(E) = {u ∈ C_{0}(R^{n}) : u|_{F} ∈ F, ∫_{R_{n}} |∇u(x)|^{2}dx < ∞},
Ẽ(u, v) = E(u|_{F} , v|_{F}) + \frac{1}{2} ∫_{R^{n}} ∇u(x)∇v(x)dx  ∀u, v ∈ D(E),

instead of the original definition given in page 71, line (-4)-(-3). Indeed, with the original Ẽ, (Ẽ, F′) is not necessarily a closed form. page 74 line 11-13 This is wrong. We should delete this sentence.